Write an exampale for a bounded sequence which is not convergent
1
Expert's answer
2021-08-05T08:07:09-0400

Let us consider "a_n=(-1)^n" . This sequence is bounded: "-1\leq a_n\leq 1" .


Suppose that it converges: "\lim\limits_{n\rightarrow \infty}a_n=A" .

Let "\varepsilon =1" . Then "\exists N" such that "\forall n>N" : "|a_n-A|<1" .

For "n=2N" we have "|1-A|<1" .

For "n=2N+1" we have "|-1-A|=|1+A|<1" .


And we have that "2>|1-A|+|1+A|\geq |1-A+1+A|=2" . So, "2>2" .


It proves that "\{a_n\}" is not convergent.